A researcher reports that mice live an average of 40 months when their diets are sharply restricted and enriched with vitamins and proteins. Assume lifetimes of such mice are normally distributed with standard deviation 6.3 months.

1. Find the proportion of mice that live more than 55 months.
1
Expert's answer
2020-09-09T18:49:25-0400

"P(X>55)=P(Z>\frac{55-40}{6.3})=P(Z>2.38)=1-P(Z<2.38)=0.0087."